# The impact of Solar magnetic field configurations on the production of gamma rays at the Solar disk

### v1

@@ 3.1 Production height of gamma rays
Figure 5: Production height of the gamma rays over the surface in different magnetic field configurations. The column denotes the magnetic field and the row the initial particle energy. The dotted line shows the minimal resolvable gamma ray density with 1 photon per spatial bin.
In Fig. 5, the distance of the produced gamma rays to the solar surface is shown. The distribution is shown for all directional filters discussed in Sec. 2.4. For a better comparison, the radial density profile is shown as a red dashed line. The minimal resolvable statistic with one observed gamma ray in a radial bin is shown as the grey dotted line.
The distribution of all produced gamma rays follows directly from the gas distribution. Only differences in the total number of photons depending on the primary energy can be observed. Two different effects lead to the change in the absolute normalization. On the one hand, the inelastic proton-proton cross section is energy dependent. This effect is important for the 1 GeV particles, as their energy is close to the kinematic threshold of the interaction. At higher energies, the cross section increases only slowly (Doerner25). On the other hand, the gamma-ray multiplicity increases for higher energies, leading to more photons in total.
@@
3.2 Initial pitch-angle distribution
The distance that a particle can travel in an increasing magnetic field before it is mirrored depends on the initial pitch angle. This effect has been analyzed for GCRs propagating from the orbit of the Earth to the outer boundary of the solar atmosphere by Hutchinson-etal-2022. Particles with a nearly perpendicular initial pitch angle will be mirrored earlier and cannot reach the lower layers of the chromosphere, where gamma rays can be produced.
@@ 3.2 Initial pitch-angle distribution
3.3 Distribution of the gamma ray production on the solar surface
The most direct impact of the magnetic field structure on gamma-ray observables is the position of produced gamma-rays projected onto the solar surface. We calculate the production position in spherical coordinates and histogram the positions with an HEALPix schema (Healpix).
In Fig. 7, the place of the gamma-ray production is shown for all produced gamma-rays. The lowest energies (left column) show only very localized places at which the gamma-ray can be produced. Only the DQCS model predicts one more pronounced band in the southern hemisphere. It should be noted that the DQCS model has cylindrical symmetry. Therefore, no longitudinal effects are expected. Also, the total number of events per pixel is quite low. A better description of the positional features would require simulations with higher statistics, which we leave for a dedicated study.
@@ 3.3 Distribution of the gamma ray production on the solar surface
Figure 8: Same as Fig. 7 for the outgoing photons. Due to the smaller statistics, the resolution is reduced to nSide=32.
@@
In Fig. 8 and Fig. 9, the maps of the production places are shown for the outgoing and passing event selection (see Sec. 2.4 for explanation). Both selections show the same spatial distribution. Therefore, we discuss these maps together and show them for completeness. In all cases, the total statistic is much lower. Therefore, the maps are converted to a lower angular resolution (nSide=32).
@@ 4 Realistic injection
4.1 Direction dependent spectrum
In the first step we shows the resulting Spectral Energy Distribution (SED) based on the different event selections discussed in Sec. 2.4 and magnetic field configurations in Fig. 10. While the PFSS models for different Carrington Rotations show a similar behavior, the energy spectrum predicted in the DQCS magnetic field configuration deviates strongly. The selection of all created gamma-rays and all gamma-rays that pass the surface does not show any break in the SED, and the spectral slope relates directly to the incoming proton SED. Only the selection of outgoing gamma-rays, which require a magnetic mirroring of the parent proton, shows a change in the spectral slope at $\sim 10\,\mathrm{GeV}$.
@@ 4.2 Data comparison
Figure 11: Spectral energy distribution of the passing gamma-rays (dash-dotted lines) compared to the observations with Fermi (Tang-etal-2018; Linden-etal-2022) and HAWC (Albert-etal-2023). The solid lines show the same simulations scaled by a factor of 12 to include missing contributions for heavier nuclei, the interplanetary magnetic field and the structures in the gas distribution (see Sec. 4.2).
@@
    GCRs do not only consist of protons but also heavier elements. Castro2025 discusses the impact of heavier elements on the total diffuse gamma-ray flux. They conclude that the total flux can be increased by a factor of up to 2 compared to a proton-only model.
@@
    For the initial pitch-angle distribution, we assume isotropic momentum at $r=2.5R_{\odot}$. This is only valid for strong scattering of particles reaching this outer boundary of the solar atmosphere. Hutchinson-etal-2022 introduces the effect of magnetic mirroring that can already happen in the Parker magnetic field between the Earth and Sun. Without additional scattering, this will change the pitch-angle distribution. In Appendix A we investigate the effect of a different pitch-angle distribution leading to an increased gamma-ray flux up to a factor 2. Depending on the details of the scattering, i.e., the pitch-angle diffusion coefficient, this effect might become even stronger.
@@
    In this study, we focus on the impact of solar magnetic field configuration. Therefore, we apply a simplified spherical symmetric gas distribution. In nature, one would expect a coupling between the magnetic field structure and the gas distribution. This will increase the total gamma-ray emission as the particles will see more column density and the interactions become more likely (Linker-etal-1999; vanDerHolst-etal-2014; Perri-etal-2022).
    Besides the increased interactions within the chromosphere, the GCRs reaching the lower boundary of our simulation, the photosphere, can interact in the deeper layers of the Sun (Li-etal-2024). The combination of both effects can increase the total gamma-ray flux easily by a factor 3 or more.
To account for the additional flux, we scale the predictions from our model in Fig. 11 by a factor of 12 and achieve a reasonable agreement with the measurements. At the lowest gamma-ray energies, the influence of the solar modulation is needed to avoid an overshooting of the Fermi measurements.
4.3 Expected Neutrino signal
Hadronic interactions do not only produce high-energy gamma rays, but also create neutrinos as a decay product of charged pions. In contrast to the gamma rays, neutrinos are weakly interacting and can therefore traverse the Sun. In a realistic description, the creation and propagation of neutrinos is necessary. Up to now, no evidence for high-energy neutrinos originating from the Sun has been found (IceCube-2025). Therefore, we start here with a simplified estimate of the expected neutrino flux from the Sun.
To convert our simulated gamma-ray flux into a neutrino signal, we follow the so-called delta-approximation (see Becker-Tjus-Merten-2020, and references therein). Positive, negative, and uncharged pions are created in equal manner, leading to a ratio of 1:3 between gamma rays and neutrinos. Here, we assume that all neutrino species have the same abundance at Earth after oscillation. The average energy of a produced neutrino is half the average energy of a gamma ray. Therefore, the total expected neutrino flux can be calculated as
@@ 4.3 Expected Neutrino signal
The flux of the neutrinos reaching the Earth can be divided into two components. The first component is connected to the gamma-ray production in the outgoing or passing direction. Those neutrinos are directed away from the Sun, and no absorption effect has to be taken into account. The second contribution are the behind the limb events. Here, the GCR interaction happens in the direction of the solar disc. While the corresponding gamma-rays are absorbed and can not be observed, some of the neutrinos can still travel through the Sun. To calculate the fraction of neutrinos passing the Sun, we assume the neutrino nucleon cross section presented by Formaggio-Zeller-2012 and a homogeneously filled sphere.
@@
The resulting estimate for the muon neutrino flux is shown in Fig. 12 together with the upper limit from the IceCube neutrino observatory (IceCube-2025). The total neutrino flux is dominated by the behind the limb events. The spectral shape is similar to the model-based upper limit derived by IceCube. The normalization is a factor $\sim 5$ lower than the limit. This is a comparable value to the IceCube results, where the limit is $2.48$ times higher than the tested model. Overall, a neutrino detection seems possible in the near future with more data and an improved detector for the IceCube upgrade.
@@
Figure 12: Expected muon neutrino flux from the hadronic interactions in the solar atmosphere. The IceCube upper limit is taken from IceCube-2025.
5 Discussion
@@
Accurate modeling of the hadronic solar gamma-ray emission with detailed spectral and spatial information remains a challenging task. Our simulations demonstrate that the large-scale solar magnetic field exerts a significant influence on both the spectral properties and the morphology of gamma-ray emission from the solar disk. The dependence of the emission on pitch-angle distributions is clearly demonstrated within our model framework. While our baseline assumption of isotropy at 2.5 $R_{\odot}$ provides a useful reference case, it is clear that the Parker spiral modifies the distribution of incoming cosmic rays well before they reach the near-Sun region. Magnetic mirroring in the heliospheric field between 1 AU and a few solar radii has been shown to preferentially select forward-directed particles, thereby potentially modifying the gamma-ray yield by allowing more of the cosmic rays with such a pitch-angle distribution to penetrate to deeper layers of the chromosphere (Hutchinson-etal-2022). This effect, together with uncertainties in the scattering conditions of the inner heliosphere, suggests that transport physics remains a key open problem for interpreting solar gamma-ray observations.
@@
Recent studies have emphasized the role of small-scale magnetic structures at the solar surface in shaping gamma-ray emission. Li-etal-2024; Li-etal-2025 demonstrated that flux-tubes and intergranular sheets can significantly increase the reflection probability of GCRs, thereby boosting the gamma-ray yield. Our results similarly support the significant impact of localized structuring, especially near active region boundaries and coronal holes. The discrepancy between our predictions and observations thus points to the need for multi-scale models that combine global coronal fields with chromospheric and photospheric fine structure, especially in target density variations. Future observations, i.e., in the multi TeV range, will help to distinguish between different production locations. While Li-etal-2025 predict an exponential cut-off at a few TeV for the gamma rays originating in the internetwork regions, the power-law in our model expands towards the $\sim 10$ TeV region.
@@
Solar modulation of the GCR spectra arriving near the Sun should only be of relevance for relatively low energies. For GeV range GCRs and solar gamma rays, it may be of interest to investigate the influence of the modulation effect on resulting gamma-ray spectra and a potential temporal signature related to the solar cycle (e.g., Acharyya-etal-2025). Additionally, transient signatures in the GeV range are also of significant interest due to Fermi observations in recent years (Ackermann-etal-2014; Ajello-2021). The primary ions causing the gamma-ray emission could originate directly from solar flare eruptions (e.g., Ryan-Lee-1991; Ryan-2000; Hurford-etal-2006; deNolfo-etal-2019; Pesce-Rollins-2024; Bruno-etal-2025), coronal mass ejection shock waves (Cliver-etal-1993; Pesce-Rollins-etal-2015; Ackermann-etal-2017; Jin-2018; Petrosian-2018; Gopalswamy-etal-2018; Kahler-etal-2018; Makela-etal-2023) or additional transient effects such as the coronal shock wave during eruptions (Warmuth-2025). In all such scenarios, the transport effects from the origin of the primary particles to the gamma-ray emission site and the associated field geometries remain a crucial factor. Our CRPropa-based model presented here is well equipped to address these questions in future studies with comprehensive end-to-end modelling.
@@
The coronal field models employed in this study, namely the PFSS model, while relatively simple to determine from magnetic maps, are limited in their capabilities to accurately describe the smaller scale structures of the solar field. While improvements to the PFSS field approach are possible (e.g. by modifying to an elliptical source surface Kruse-2021), more sophisticated models based on non-linear force-free extrapolations (Wiegelmann-2004) or inputs directly from MHD simulations (Jin-etal-2016) should be studied in the future.

### v2

@@ III.1 Production height of gamma rays
Figure 5: Production height of the gamma rays over the surface in different magnetic field configurations. The row denotes the magnetic field configuration and the column the initial particle energy. The dotted line shows the minimal resolvable gamma ray density with 1 photon per spatial bin.
In Fig. 5, the distance of the produced gamma rays to the solar surface is shown. The distribution is shown for all directional filters discussed in Sec. II.4. For a better comparison, the radial density profile is shown as a red dashed line. The minimal resolvable statistic with one observed gamma ray in a radial bin is shown as the gray dotted line.
The distribution of all produced gamma rays follows directly from the gas distribution. Only differences in the total number of photons depending on the primary energy can be observed. Two different effects lead to the change in the absolute normalization assuming the same column density. On the one hand, the inelastic proton-proton cross section is energy dependent. This effect is important for the 1 GeV particles, as their energy is close to the kinematic threshold of the interaction. At higher energies, the cross section increases only slowly (Dörner et al., 2025). On the other hand, the gamma-ray multiplicity increases for higher energies, leading to more photons in total. Changes in the absolute normalization due to solar modulation have not been taken into account for this simulation.
@@
III.2 Trapping of cosmic rays
The total column density, which the GCRs transverse depends on the efficiency of trapping the particles in horizontal magnetic field structures. To investigate this effect in the given field configurations, we calculate the time between the injection of GCRs at the outer boundary and the interaction in the solar atmosphere.
In Fig. 6 the emission rate of secondary gamma-rays is shown. In all magnetic field configurations clear differences between the lower ($E=1$ GeV or $E=10^{2}$ GeV) and high ($E=10^{4}$ GeV) can be seen. Only GCRs with lower energy can enter loops of closed field lines and are efficiently trapped in the horizontal field. This increases the total confinement significantly and increases the production rate at later times ($t>10^{2}$ s).
At the highest energy the difference in the field geometry between the more ordered solar minimum cases (DQCS or CR 2224) and the chaotic solar maximum cases (CR 2154 or 2157) can be seen. The more irregular structure at the solar maximum allows even the high energetic GCRs to enter closed field lines. This leads to slightly higher confinement times, but is still less than in the lower energetic case.
The horizontal structure of these loops can explain the difference between the absolute normalization in the all and outgoing/passing event selection as seen in Fig. 5.
Figure 6: Production rate of secondary gamma-rays dependent on the propagation time of their parent GCRs. The time is taken from their starting surface at $r=2.5R_{\odot}$ until interacting in the solar atmosphere. Here, all produced gamma-rays are shown.
III.3 Initial pitch-angle distribution
The distance that a particle can travel in an increasing magnetic field before it is mirrored depends on the initial pitch angle. This effect has been analyzed for GCRs propagating from the orbit of the Earth to the outer boundary of the solar atmosphere by Hutchinson et al. (2022). Particles with a nearly perpendicular initial pitch angle will be mirrored earlier and cannot reach the lower layers of the chromosphere, where gamma rays can be produced.
@@ III.3 Initial pitch-angle distribution
III.4 Distribution of the gamma ray production on the solar surface
The most direct impact of the magnetic field structure on gamma-ray observables is the position of produced gamma-rays projected onto the solar surface. We calculate the production position in spherical coordinates and histogram the positions with an HEALPix schema (Górski et al., 2005).
In Fig. 8, the place of the gamma-ray production is shown for all produced gamma-rays. The lowest energies (left column) show only very localized places at which the gamma-ray can be produced. Only the DQCS model predicts one more pronounced band in the southern hemisphere. It should be noted that the DQCS model has cylindrical symmetry. Therefore, no longitudinal effects are expected. These latitudinal stripes correspond to the open field originating from the intersection of the quadrupole loops (compare Fig. 2). Also, the total number of events per pixel is quite low. A better description of the positional features would require simulations with higher statistics, which we leave for a dedicated study.
@@ III.4 Distribution of the gamma ray production on the solar surface
Figure 9: Same as Fig. 8 for the outgoing photons. Due to the smaller statistics, the resolution is reduced to nSide=32 corresponding to a 16 times larger pixel size.
@@
In Fig. 9 and Fig. 10, the maps of the production places are shown for the outgoing and passing event selection (see Sec. II.4 for explanation). Both selections show the same spatial distribution. Therefore, we discuss these maps together and show them for completeness. In all cases, the total statistic is much lower. Therefore, the maps are converted to a lower angular resolution (nSide=32).
@@
With current instruments an observation of the two dimensional distribution of the gamma-rays on the solar surface is not feasible. Therefore, we compute latitudinal profiles similar to the analysis presented in Linden et al. (2018). Here, we restrict the analysis only to the surface maps from the passing event selection displayed in Fig. 10, as this is the most realistic one to be compared to observation. The resulting profiles are shown in Fig. 11. Our simulation are performed for mono-energetic GCR protons and consider photons of all energies. Therefore, a direct comparison to the observation by Linden et al. (2018) is not possible. Possible tensions between our prediction and the observations, have to been investigated further using more realistic magnetic field models and higher statistics observation.
Figure 11: Latitudinal profiles of the gamma-ray distribution calculated from the map of passing photons shown in Fig. 10.
@@ IV Realistic injection
IV.1 Direction dependent spectrum
In the first step we shows the resulting Spectral Energy Distribution (SED) based on the different event selections discussed in Sec. II.4 and magnetic field configurations in Fig. 12. While the PFSS models for different Carrington Rotations show a similar behavior, the energy spectrum predicted in the DQCS magnetic field configuration deviates strongly. The selection of all created gamma-rays and all gamma-rays that pass the surface does not show any break in the SED, and the spectral slope relates directly to the incoming proton SED. Only the selection of outgoing gamma-rays, which require a magnetic mirroring of the parent proton, shows a change in the spectral slope at $\sim 10\,\mathrm{GeV}$.
@@ IV.2 Data comparison
Figure 13: Spectral energy distribution of the passing gamma-rays (dash-dotted lines) compared to the observations with Fermi (Tang et al., 2018; Linden et al., 2022) and HAWC (Albert et al., 2023). The solid lines show the same simulations scaled by a factor of 12 to include missing contributions for heavier nuclei, the interplanetary magnetic field and the structures in the gas distribution (see Sec. IV.2).
@@
    GCRs do not only consist of protons but also heavier elements. Castro et al. (2025) discusses the impact of heavier elements on the total diffuse gamma-ray flux. These authors use measurements for the contribution by GCR proton and helium and estimate limits for a heavier component by assuming a carbon or iron dominance. They conclude that the total flux can be increased by a factor of up to 2 compared to a proton-only model.
@@
    For the initial pitch-angle distribution, we assume isotropic momentum at $r=2.5R_{\odot}$. This is only valid for strong scattering of particles reaching this outer boundary of the solar atmosphere. Hutchinson et al. (2022) introduces the effect of magnetic mirroring that can already happen in the Parker magnetic field between the Earth and Sun. Without additional scattering, this will change the pitch-angle distribution. In Appendix A we investigate the effect of a different pitch-angle distribution leading to an increased gamma-ray flux up to a factor 2. Depending on the details of the scattering, i.e., the pitch-angle diffusion coefficient, this effect might become even stronger.
@@
    In this study, we focus on the impact of solar magnetic field configuration. Therefore, we apply a simplified spherical symmetric gas distribution. In nature, one would expect a coupling between the magnetic field structure and the gas distribution. State-of-the-art 3D MHD models, by solving the coupled plasma–magnetic system, can provide physically self-consistent 3D densities that vary with magnetic topology and dynamics, and may yield higher densities in specific regions than a 1D model. This will increase the total gamma-ray emission as the particles will see more column density and the interactions become more likely (Linker et al., 1999; van der Holst et al., 2014; Perri et al., 2022).
    Besides the increased interactions within the chromosphere, the GCRs reaching the lower boundary of our simulation, the photosphere, can interact in the deeper layers of the Sun (Li et al., 2024a). The combination of both effects can increase the total gamma-ray flux easily by a factor 3 or more.
To account for the additional flux, we scale the predictions from our model in Fig. 13 by a factor of 12 and achieve a reasonable agreement with the measurements. At the lowest gamma-ray energies, the influence of the solar modulation, which is not included in the model presented here, is needed to avoid an overshooting of the Fermi measurements.
IV.3 Expected Neutrino signal
Hadronic interactions do not only produce high-energy gamma rays, but also create neutrinos as a decay product of charged pions. In contrast to the gamma rays, neutrinos are weakly interacting and can therefore traverse the Sun. In a realistic description, the creation and propagation of neutrinos is necessary. Up to now, no evidence for high-energy neutrinos originating from the Sun has been found (Abbasi et al., 2025). Therefore, we start here with a simplified estimate of the expected neutrino flux from the Sun.
To convert our simulated gamma-ray flux into a neutrino signal, we follow the so-called delta-approximation (see Becker Tjus and Merten, 2020, and references therein). Positive, negative, and uncharged pions are created in equal manner, leading to a ratio of 1:3 between gamma rays and neutrinos. Here, we assume that all neutrino species have the same abundance at Earth after oscillation. The average energy of a produced neutrino is half the average energy of a gamma ray. Therefore, the total expected neutrino flux can be calculated as
@@ IV.3 Expected Neutrino signal
The flux of the neutrinos reaching the Earth can be divided into two components. The first component is connected to the gamma-ray production in the outgoing or passing direction. Those neutrinos are directed away from the Sun, and no absorption effect has to be taken into account. The second contribution are the behind the limb events. Here, the GCR interaction happens in the direction of the solar disc. While the corresponding gamma-rays are absorbed and can not be observed, some of the neutrinos can still travel through the Sun. To calculate the fraction of neutrinos passing the Sun, we assume the neutrino nucleon cross section presented by Formaggio and Zeller (2012) and a homogeneously filled sphere.
@@
The resulting estimate for the muon neutrino flux is shown in Fig. 14 together with the upper limit from the IceCube neutrino observatory (Abbasi et al., 2025). The total neutrino flux is dominated by the behind the limb events. The spectral shape is similar to the model-based upper limit derived by IceCube. The normalization is a factor $\sim 5$ lower than the limit. This is a comparable value to the IceCube results, where the limit is $2.48$ times higher than the tested model. Overall, a neutrino detection seems possible in the near future with more data and an improved detector for the IceCube upgrade.
@@
Figure 14: Expected muon neutrino flux from the hadronic interactions in the solar atmosphere. The IceCube upper limit is taken from Abbasi et al. (2025).
V Discussion
@@
Accurate modeling of the hadronic solar gamma-ray emission with detailed spectral and spatial information remains a challenging task. Our simulations demonstrate that the large-scale solar magnetic field exerts a significant influence on both the spectral properties and the morphology of gamma-ray emission from the solar disk. The dependence of the emission on pitch-angle distributions is clearly demonstrated within our model framework. While our baseline assumption of isotropy at 2.5 $R_{\odot}$ provides a useful reference case, it is clear that the Parker spiral modifies the distribution of incoming cosmic rays well before they reach the near-Sun region. Magnetic mirroring in the heliospheric field between 1 AU and a few solar radii has been shown to preferentially select forward-directed particles, thereby potentially modifying the gamma-ray yield by allowing more of the cosmic rays with such a pitch-angle distribution to penetrate to deeper layers of the chromosphere (Hutchinson et al., 2022). This effect, together with uncertainties in the scattering conditions of the inner heliosphere, suggests that transport physics remains a key open problem for interpreting solar gamma-ray observations.
@@
Recent studies have emphasized the role of small-scale magnetic structures at the solar surface in shaping gamma-ray emission. Li et al. (2024a, 2025) demonstrated that flux-tubes and intergranular sheets can significantly increase the reflection probability of GCRs, thereby boosting the gamma-ray yield. Our results similarly support the significant impact of localized structuring, especially near active region boundaries and coronal holes. The discrepancy between our predictions and observations thus points to the need for multi-scale models that combine global coronal fields with chromospheric and photospheric fine structure, especially in target density variations. Future observations, i.e., in the multi TeV range, will help to distinguish between different production locations. While Li et al. (2025) predict an exponential cut-off at a few TeV for the gamma rays originating in the internetwork regions, the power-law in our model expands towards the $\sim 10$ TeV region.
@@
Solar modulation of the GCR spectra arriving near the Sun should only be of relevance for relatively low energies. For GeV range GCRs and solar gamma rays, it may be of interest to investigate the influence of the modulation effect on resulting gamma-ray spectra and a potential temporal signature related to the solar cycle (e.g., Acharyya et al., 2025). Additionally, transient signatures in the GeV range are also of significant interest due to Fermi observations in recent years (Ackermann et al., 2014; Ajello et al., 2021). The primary ions causing the gamma-ray emission could originate directly from solar flare eruptions (e.g., Ryan and Lee, 1991; Ryan, 2000; Hurford et al., 2006; de Nolfo et al., 2019; Pesce-Rollins et al., 2024; Bruno et al., 2025), coronal mass ejection shock waves (Cliver et al., 1993; Pesce-Rollins et al., 2015; Ackermann et al., 2017; Jin et al., 2018; Petrosian, 2018; Gopalswamy et al., 2018; Kahler et al., 2018; Mäkelä et al., 2023) or additional transient effects such as the coronal shock wave during eruptions (Warmuth et al., 2025). In all such scenarios, the transport effects from the origin of the primary particles to the gamma-ray emission site and the associated field geometries remain a crucial factor. Our CRPropa-based model presented here is well equipped to address these questions in future studies with comprehensive end-to-end modelling.
@@
The coronal field models employed in this study, namely the PFSS model, while relatively simple to determine from magnetic maps, are limited in their capabilities to accurately describe the smaller scale structures of the solar field. While improvements to the PFSS field approach are possible (e.g. by modifying to an elliptical source surface Kruse et al., 2021), more sophisticated models based on non-linear force-free extrapolations (Wiegelmann, 2004) or inputs directly from MHD simulations (Jin et al., 2016) should be studied in the future.